OpenStudy (anonymous):

Check please! c: A line passes through (–5, 8) and (3, 2). What is the slope of the line? A. -3/4 <<<< B. -4/3 C. 3 D. 1/3

OpenStudy (ineedhelpfast12):

B

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

I couldv'e sworn I got it right this time. :C

OpenStudy (ineedhelpfast12):

slope is rise over run, so y over x. x would be the distance from -5 to 3, which equals 8.

OpenStudy (ineedhelpfast12):

then x would be the distance from 8 to 2, which is -6, so 8 over 6 simplified is 4/3, then add the negative from the 6, and the answer will be -4/3

OpenStudy (anonymous):

So what did I do wrong?

jimthompson5910 (jim_thompson5910):

@Ineedhelpfast12 you mixed up x and y

OpenStudy (anonymous):

So did I do it right? @jim_thompson5910

OpenStudy (ineedhelpfast12):

whopsiedaisy

OpenStudy (anonymous):

yes

OpenStudy (ineedhelpfast12):

no what I posted is right, the first post was supposed to say y instead of x

jimthompson5910 (jim_thompson5910):

slope = (y2-y1)/(x2-x1) slope = (2-8)/(3 - (-5)) slope = (-6)/(8) slope = -3/4

OpenStudy (anonymous):

eight to the right, six down \[-\frac{6}{8}=-\frac{3}{4}\]

OpenStudy (ineedhelpfast12):

@jim_thompson5910 4 is the rise, 3 is the run. to write a number in slope intercept form, it has to be rise over run, so -4/3

OpenStudy (ineedhelpfast12):

@satellite73 it is 8 up, 6 to the right